• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Templon

    Log entry time 10:37:47 on December 18,1999

    Entry number 31394

    keyword=cryotarget, IOC reboot, temp weirdness

    cryotarget -- noted temperatures significantly (0.15 deg) above setpt.
    Associated with raising beam current. Heater still running (ie had
    more headroom to cool target) -- called Kuss. rebooted IOC (he
    could not connect to some devices from home).

    On further inspection, find now that (after reboot) temperatures are
    still higher than desired on CT171 and CT168 (the TV temps) but
    other CT's (esp. one on which we are feeding back, CT 97) are all
    fine at 6.3 K. Thus for some reason the two CT's closest to the can
    are reading 0.15 K higher than all the ones before. Possible candidates:
    deterioration of these temperature sensors, or fan in initial stages of
    failure. Or of course, something else.